OpenStudy (anonymous):

I have three numbers. The biggest one is twice the middle one, and the biggest one plus the middle one is four times the smallest one. The smallest one plus the middle one is two less than the biggest one. What are the numbers?

OpenStudy (anonymous):

B = biggest number M = middle number S = smallest number B = 2M B+M=4S S+M=B-2 plug in B=2M for both equations: 2M+M = 4S S+M = 2M-2, OR: M = S+2 Plug in M = S+2: 3(S+2) = 4S 3S + 6 = 4S S = 6 3M = 4(6) M = 8 B=2(8)=16 Smallest number is 6 Medium number is 8 Largest is 16

OpenStudy (anonymous):

wow